# A Random Rule Model

### 04 Mar 2026 v1

@@ 4.1 Rule-level Diagnostics
The pair $(\phi(f),\sigma_{N}(f))$ separates distinct roles. A large $\phi(f)$ indicates that the procedure improves predictive fit and is not easily replaced by the rest of the library; the sign and magnitude of $\sigma_{N}(f)$ indicate whether that procedure also shapes the overall mixture structure (concentrating or diversifying effective responsibility). In the Results section we report $(w_{f},\phi(f),\sigma_{N}(f))$ side-by-side to identify rules that are heavily used, rules that are essential for accuracy, and rules that mainly act as substitutes or robustness variants.
5 Benchmark models
$\hat{p}(A)=\frac{\exp(\eta V(L^{1}))}{\exp(\eta V(L^{1}))+\exp(\eta V(L^{2}))},$ (18)
where $\eta$ is a free “determinism” (inverse-noise) parameter. The exception is the fully context-dependent network, which directly outputs $\hat{p}(A)$.
Neural EU.
Neural PT.
Neural CPT.
Value-Based.
$V(L)=f(x^{L},p^{L}),$
Context-Dependent.
$\hat{p}(A)=g(x^{L^{1}},p^{L^{1}},x^{L^{2}},p^{L^{2}}),$
Mixture of Theories (MOT).
6 Empirical design
This section explains what is being predicted, how menus are encoded, how each model is trained. The core object is the menu-level choice probability $p(A)=\Pr(L^{1}\mid A)$; every model outputs a prediction $g(A)\in[0,1]$ that we evaluate out of sample.
@@ 6.1 Dataset and target variable
We use the choices13k dataset introduced by Peterson et al. (2021), containing 13,006 binary risky-choice problems. Each problem corresponds to a menu $A=\{L^{1},L^{2}\}$ and is repeated across many observations, allowing the computation of a menu-level choice frequency. We work with the filtered dataset that (i) removes ambiguous problems (masked probabilities) and (ii) keeps only problems for which feedback was provided, yielding N=9,831N=9{,}831 menus. The feedback restriction follows the baseline protocol of Peterson et al. (2021) and has a practical justification: feedback problems have more repetitions per menu, producing more precise estimates of the choice frequency $\hat{p}(A)$ that serves as our prediction target. Because our model is static (it maps menu features to a predicted choice probability without using feedback information), the feedback condition affects only the precision of the dependent variable, not the model’s information set. [^5] Feedback can in principle induce learning effects. If such dynamics are present, our static model captures the average behavior under the feedback regime. For each menu $A$, the dataset reports the fraction of times the left option is chosen. We denote this empirical frequency by $\hat{p}(A)$ and treat it as the target to be predicted.
@@ 6.2 Evaluation metric
Because the number of trials $n_{t}$ behind each menu-level frequency $\hat{p}(A_{t})$ varies across menus, we also report a trial-weighted variant:
$\mathrm{MSE}_{w}(g)=\frac{\sum_{t=1}^{T}n_{t}\Big(g(A_{t})-\hat{p}(A_{t})\Big)^{2}}{\sum_{t=1}^{T}n_{t}}.$ (20)
@@ 6.4 Cross-validation protocol
6.5 Estimation of the rule-gating model
The rule-gating model differs from the benchmark neural predictors in a key way: the rules are deterministic and parameter-free. Estimation concerns only the gate, i.e. the mapping from menu features $z(A)$ to rule propensities $q(A)\in\Delta^{F-1}$ (see Section 6.6 for the definition of $z(A)$).
@@ 6.5 Estimation of the rule-gating model
Interpretation.
The estimated gate $q(A;\theta)$ can be read as a context-dependent mixture over simple procedures: for menus where a rule is inactive, it is effectively excluded by the activity filter; among active rules, the gate distributes probability mass so that the resulting mixture best matches observed choice frequencies.
@@ 6.6 Gate features
The gate’s role is to modulate rule responsibilities using observable features of the menu. A natural question is which features to use. We adopt as the baseline specification a $12$-dimensional vector of economically motivated summary statistics $z(A)\in\mathbb{R}^{12}$. Six are pairwise gaps between the two lotteries: the expected-value gap $\mathrm{EV}(L^{1})-\mathrm{EV}(L^{2})$, the best-case outcome gap $\max(L^{1})-\max(L^{2})$, the worst-case outcome gap $\min(L^{1})-\min(L^{2})$, the variance gap $\mathrm{Var}(L^{1})-\mathrm{Var}(L^{2})$, the modal-outcome gap $\mathrm{Mode}(L^{1})-\mathrm{Mode}(L^{2})$, and the skewness gap $\mathrm{Skew}(L^{1})-\mathrm{Skew}(L^{2})$. These gaps are designed to capture the contrasts that the rules in the library operate on: the min/max gaps relate to MMn/MMx, the mode gap relates to MAP, and the EV and variance gaps relate to risk-return tradeoffs. The remaining six are level variables: the expected values $\mathrm{EV}(L^{1})$ and $\mathrm{EV}(L^{2})$, the standard deviations $\mathrm{SD}(L^{1})$ and $\mathrm{SD}(L^{2})$, the largest absolute outcome across both lotteries $\max_{i\in\{1,2\}}\max|\mathrm{Supp}(L^{i})|$, and the support-size gap $|\mathrm{Supp}(L^{1})|-|\mathrm{Supp}(L^{2})|$. The level variables allow the gate to condition rule propensities on the scale and complexity of the menu environment. By construction, $d_{z}=12$ is much smaller than the $4M=40$ dimensions of the raw encoding, but it captures the economically relevant variation: each feature has a direct verbal interpretation as a lottery comparison or a menu characteristic.
@@
7 Results
This section reports out-of-sample predictive performance of the learned random rule model relative to the benchmark suite of Peterson et al. (2021), followed by completeness and restrictiveness diagnostics and a portability evaluation.
7.1 Predictive performance
Table 1 reports mean test Mean Square Error (MSE) across 50 random train/test splits, with learning-rate selection as described in Section 6. The main finding is that the rule-gating predictor substantially outperforms the standard structured models (Neural EU/PT/CPT and Value-Based), while remaining competitive with the two strongest flexible benchmarks.
Concretely, rule-gating improves on the Value-Based model and on all Neural EU/PT/CPT baselines by a substantial margin. These gains indicate that a small repertoire of transparent “lines of conduct,” combined with a flexible gate that learns when each procedure applies, captures substantial regularity in the data that is missed by standard preference-based architectures. Rule-gating also outperforms the flexible Context-Dependent model. Only the Mixture of Theories (MOT) achieves a lower MSE, consistent with the fact that this model allows richer menu interactions than any fixed rule library. [^6] A caveat on feature sets: the rule-gating model’s effective inputs include both the $12$-dimensional gate features $z(A)$ and the $24$-dimensional rule activity and recommendation indicators, which are nonlinear functions of the lottery primitives. The neural benchmarks operate on the raw $40$-dimensional encoding $\psi(A)$ with small, generic architectures. The comparison therefore does not hold features fixed; the sklearn comparison in the appendix provides the closest feature-controlled evaluation.
Figure 1(a) visualizes the same comparison with uncertainty bands across splits. Performance differences are stable across splits, and the ranking is preserved across learning-rate choices in our grid. Figure 1(b) reports a complementary “learning-curve” diagnostic, plotting test-set MSE as a function of the fraction of training data used. Two patterns emerge. First, the learning curve of rule-gating tracks that of the fully context-dependent network and outperforms it across training fractions, suggesting that an interpretable rule-based structure extracts at least as much predictive content as a fully flexible context-dependent architecture. Second, rule-gating exhibits a clear and persistent improvement over the value-based network throughout the range of training fractions, consistent with the view that allowing menu-dependent selection among a small repertoire of procedures adds predictive content beyond a single flexible value representation. Finally, MOT continues to perform best across the curve, reflecting the additional flexibility afforded by a richer mixture-of-experts structure.
The trial-weighted metric $\mathrm{MSE}_{w}$ (Table 1) confirms that the ranking of models is stable when predictions are weighted by the precision of the observed choice frequencies: all pairwise orderings are preserved, and the magnitude of the gaps changes only marginally. Appendix Table A.2 reports split-level comparisons: for each pair of models, per-split metric differences are computed at the best learning rate, and we report the mean and standard error across 50 splits. Appendix Table A.3 reports split-variability intervals constructed from fold-to-fold variability across the 50 random splits. These intervals summarize how performance varies with the resampling of train/test partitions.
Interpretability–accuracy tradeoff.
A gradient-boosted tree ensemble (GBT) trained on raw lottery summary statistics achieves MSE $0.0100$ (Appendix Table A.9), meaningfully below both rule-gating ($0.0128$) and MOT ($0.0114$). The interpretability cost of the RRM is therefore non-trivial. We view this as a worthwhile tradeoff because the RRM’s interpretability is structural, not post-hoc: for each menu, it decomposes the prediction into contributions from named behavioral procedures with economically interpretable weights. Post-hoc explainability tools identify which features matter but cannot decompose predictions into named procedures. This structural decomposition, together with the identification guarantees of Section 3, is what permits the mechanism-level analysis of Section 8.
Model Mean MSE SD Mean $\mathrm{MSE}_{w}$ Best LR
MOT 0.01139 0.00056 n.a. 0.001
Rule-gating 0.01282 0.00059 0.01269 0.010
Context-dependent 0.01344 0.00081 0.01330 0.010
Value-based 0.01921 0.00081 0.01904 0.010
Neural PT 0.02064 0.00085 0.02050 0.010
Neural CPT 0.02145 0.00093 0.02134 0.010
Neural EU 0.02215 0.00085 0.02205 0.010
Note: MOT learning rate is selected by test MSE following the published protocol of Peterson et al. (2021) (see Section 6.4); the HURD framework does not produce per-menu predictions, so $\mathrm{MSE}_{w}$ is unavailable for MOT in cross-validation. Rule-gating here uses the full 12-rule library. Approximate parameter counts: Neural EU/PT/CPT $\approx 60$$120$; Value-based $\approx 200$; Context-dep. $\approx 200$; Rule-gating $143$; MOT: see Peterson et al. (2021).
(a) Mean test MSE across 50 random splits (split-variability band).
7.2 Completeness and Restrictiveness
Table 1 reports absolute out-of-sample predictive performance across models. In this section we add two complementary diagnostics that help interpret these performance differences. First, we measure completeness: how much of the predictive gains delivered by a highly flexible benchmark can be recovered by a more structured (and potentially interpretable) model. Second, we measure restrictiveness: how strongly a model class constrains the set of menu-to-choice mappings it can fit. Jointly, these measures help distinguish performance gains that reflect substantive structure from gains that could arise from sheer flexibility.
Machine Learning (ML) Completeness Score
While Table 1 reports absolute predictive performance, it is also useful to quantify how much of the predictive gains delivered by the strongest flexible architecture can be recovered by an interpretable model. To this end, we compute a ML completeness measure inspired by recent work on the completeness of economic models (e.g., Peysakhovich and Naecker (2017), Fudenberg and Puri (2022), Fudenberg and Liang (2020), and Fudenberg et al. (2022)). We take Neural EU as the baseline model, since it is the most basic model in our benchmark class, and we take the Mixture of Theories (MOT) as the flexible benchmark. For any model class $g$, we define its ML completeness score by
This statistic measures the fraction of the baseline-to-MOT improvement in out-of-sample MSE captured by $g$. By construction, $\mathrm{Comp}(g_{\mathrm{EU}})=0$ and $\mathrm{Comp}(g_{\mathrm{MOT}})=1$. Values in $(0,1)$ indicate partial recovery of MOT’s predictive gains, values above $1$ indicate that a model outperforms MOT, and negative values indicate performance below the baseline. The score depends on the choice of endpoints: using a different baseline (e.g., Neural PT instead of Neural EU) or a different flexible benchmark would rescale the numbers, though ordinal rankings across models would be preserved.
Figure 2(a) visualizes the ML completeness scores implied by Table 1. The rule-gating model closes a substantial fraction of the baseline-to-MOT performance gap ($\approx 87\%$), markedly exceeding the Value-based architecture ($\approx 27\%$). This normalization emphasizes that the predictive gains delivered by rule-gating are not marginal improvements over the expected utility benchmark, but a sizable step toward the performance of the strongest flexible models.
Restrictiveness
(a) ML Completeness Scores $\mathrm{Comp}(g)$
(b) Restrictiveness $\mathrm{Restr}(g)$
Figure 2: Completeness and Restrictiveness.
Model-wise restrictiveness.
Figure 2(b) reports $\mathrm{Restr}(g)$ for each model. MOT stands out as markedly less restrictive ($\mathrm{Restr}\approx 0.2$), and the context-dependent architecture is moderately flexible ($\mathrm{Restr}\approx 0.9$). All other models—including rule-gating and the neural utility benchmarks—have restrictiveness very close to $1$, implying that they cannot improve meaningfully over a constant predictor on permuted targets.
Figure 3 plots each model in the (completeness, restrictiveness) plane. The neural utility benchmarks cluster at high restrictiveness but low completeness; MOT achieves maximal completeness with low restrictiveness. Rule-gating and context-dependent occupy the upper-right region: the context-dependent model is more complete but less restrictive, whereas rule-gating is more restrictive but somewhat less complete. These results suggest that rule-gating delivers a large share of the predictive gains of flexible benchmarks while retaining strong discipline against fitting arbitrary mappings.
7.3 Portability
To assess portability, we evaluate the models trained on choices13k (Peterson et al., 2021) on an independent dataset of binary lottery choices, the publicly available CPC18 competition dataset (Erev et al., 2017; Plonsky et al., 2017). [^7] The data can be downloaded at https://cpc-18.com/data/. We restrict attention to the risk problems and exclude ambiguity tasks. The resulting sample comprises $N=686$ subjects recruited across multiple locations in Israel. Each subject faces $20$ to $28$ distinct binary lottery choice problems, and each problem is repeated for $25$ trials, yielding $500$ to $700$ decisions per subject. The design includes both no-feedback and feedback regimes.
Because choices13k reports aggregate choice frequencies by menu, we aggregate CPC18 to the menu level by computing, for each unique binary menu $A$, the empirical left-choice frequency $\hat{p}_{\mathrm{CPC}}(A)$ across all available repetitions (and subjects). This produces a dataset of $|\mathcal{A}^{\mathrm{CPC}}|=180$ distinct menus and 354,400354{,}400 total trials (the product of the $686$ subjects and their individual problem counts, each repeated $25$ times). [^8] In addition to the unweighted menu-level aggregation, we also consider a trial-weighted evaluation that weights each menu by its number of observations. Although CPC18 is much smaller than choices13k, it provides a stringent out-of-domain test of whether a learned mapping from menu features to choice probabilities generalizes beyond the environment on which it was trained.
Portability design.
For each model class, we estimate parameters on choices13k exactly as in Section 6.4, using the model-specific learning rate selected in Table 1. We then freeze these parameters and apply the fitted predictor to the CPC18 menus. Let $\mathcal{A}^{13k}$ denote the filtered choices13k analysis set and let $\mathcal{A}^{\mathrm{CPC}}$ denote the set of binary menus in CPC18 after applying the same basic preprocessing (e.g., removing masked probabilities when relevant). Using the parameters learned on choices13k, each model produces a portability prediction $g_{\rightarrow}(A)\in[0,1]$ for each $A\in\mathcal{A}^{\mathrm{CPC}}$. We evaluate portability using the same mean-squared error metric:
All models use the same menu encoding $\psi(A)$ as in (21). The outcome rescaling factor (maximum absolute payoff) is computed from the choices13k training sample and applied identically to CPC18 outcomes, so that the feature space is consistent across datasets. [^9] Because both datasets use a common monetary unit (points/tokens), the rescaling factor simply puts all outcomes on a common numerical scale. No currency conversion is required. No hyperparameters are tuned on CPC18: $\mathrm{MSE}_{\mathrm{CPC}}$ is a pure out-of-domain test. Models are exported without re-tuning.
Trial-level evaluation.
Because CPC18 records individual-level trial data, we also evaluate portability at the trial level using the Brier score and Bernoulli log-loss, which assess each menu-level prediction against realized binary choices. Trial-level evaluation is the primary portability metric; menu-level MSE is retained for comparability (Appendix Table A.7).
Portability results.
Table 2 reports portability metrics for the full set of models. Two patterns stand out. First, models that perform well in-sample on choices13k are not necessarily portable: notably, MOT—the most flexible benchmark and the best performer in Table 1—performs poorly under transfer, suggesting that a substantial share of its predictive advantage is dataset-specific and does not export cleanly to a different menu distribution and subject pool. Second, the rule-gating model retains strong predictive content under transfer: it achieves CPC18 MSE $0.0170$, essentially tied with the context-dependent model ($0.0172$) and far ahead of the structured neural benchmarks ($\approx 0.032$). The margin between rule-gating and the context-dependent model is small relative to the CPC18 sample size ($180$ menus) and should not be over-interpreted; the salient finding is that both models transfer well while MOT does not. Rule-gating therefore offers strong portability with an interpretable behavioral decomposition, whereas the context-dependent model — despite similar transfer accuracy — is a black-box predictor.
Rule-gating 0.1971 0.5747 0.0170
Context-dependent 0.1977 0.5760 0.0172
Neural CPT 0.2111 0.6112 0.0320
Neural EU 0.2114 0.6122 0.0320
Value-based 0.2120 0.6142 0.0321
Neural PT 0.2121 0.6136 0.0326
MOT 0.2691 0.7480 0.0926
8 Mechanisms

### 19 Mar 2026 v2

@@ 4.1 Rule-level Diagnostics
The pair $(\phi(f),\sigma_{N}(f))$ separates distinct roles. A large $\phi(f)$ indicates that the procedure improves predictive fit and is not easily replaced by the rest of the library; the sign and magnitude of $\sigma_{N}(f)$ indicate whether that procedure also shapes the overall mixture structure (concentrating or diversifying effective responsibility). In Section 7 we report $(w_{f},\phi(f),\sigma_{N}(f))$ side-by-side to identify rules that are heavily used, rules that are essential for accuracy, and rules that mainly act as substitutes or robustness variants.
5 Estimation and inference
The identification results suggest a constructive two-step estimator. The first step recovers, at feature values satisfying the fixed-feature rank condition, the vector of relative rule weights. The second step uses the affine restriction on log-weights to recover the gate parameters. This section treats the econometric object of interest—latent procedural heterogeneity—separately from the gradient-based prediction procedure used later for model comparison.
Sampling scheme.
Let $\mathcal{M}=\{A_{1},\ldots,A_{M}\}$ be a finite collection of menus. For each menu $A_{m}$, we observe $n_{m}$ independent binary choices $Y_{mi}\in\{0,1\}$, $i=1,\ldots,n_{m}$, with
$\Pr(Y_{mi}=1\mid A_{m})=p(A_{m})\in(0,1).$
For distinct menus, the menu-specific samples are mutually independent. Write
$\hat{p}_{m}:=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}Y_{mi},\qquad\hat{r}_{m}:=\frac{\hat{p}_{m}}{1-\hat{p}_{m}},$
where, for numerical stability only, $\hat{p}_{m}$ may be trimmed to $\varepsilon_{n},1-\varepsilon_{n}]$ with $\varepsilon_{n}\downarrow 0$ slowly enough that trimming is asymptotically inactive.
We study the asymptotic regime
$N:=\min_{1\leq m\leq M}n_{m}\to\infty,\qquad\frac{n_{m}}{N}\to\pi_{m}\in(0,\infty)\quad\text{for each }m.$
Identifying feature values.
Let $x^{(1)},\ldots,x^{(K)}$ be distinct feature values with $K\geq d+1$, and define the associated menu cells
$\mathcal{M}_{k}:=\{m\in\{1,\ldots,M\}:\psi(A_{m})=x^{(k)}\}.$
For each $k$, let $M_{k}:=|\mathcal{M}_{k}|$ and assume $M_{k}\geq|\mathcal{F}|-1$. For $m\in\mathcal{M}_{k}$, define
$h_{f}(A_{m}):=\kappa_{f}^{L}(A_{m})-r(A_{m})\kappa_{f}^{R}(A_{m}),\qquad\hat{h}_{f}(A_{m}):=\kappa_{f}^{L}(A_{m})-\hat{r}_{m}\,\kappa_{f}^{R}(A_{m}).$
Let $H^{(k)}$ and $\hat{H}^{(k)}$ denote the $M_{k}\times|\mathcal{F}|$ matrices stacking the row vectors $h(A_{m})^{\top}$ and $\hat{h}(A_{m})^{\top}$ over $m\in\mathcal{M}_{k}$.
Fix the baseline rule $f_{0}$, and reorder coordinates so that $f_{0}$ is first. Partition
$H^{(k)}=\big[h_{0}^{(k)}\ \ H_{-}^{(k)}\big],\qquad\hat{H}^{(k)}=\big[\hat{h}_{0}^{(k)}\ \ \hat{H}_{-}^{(k)}\big],$
where $h_{0}^{(k)}$ is the first column and $H_{-}^{(k)}$ collects the remaining $|\mathcal{F}|-1$ columns.
First-stage estimator: cellwise rule weights.
Under Theorem 1, the normalized weight vector at $x^{(k)}$ is uniquely determined by $\omega_{f_{0}}(x^{(k)})=1$ and
$H^{(k)}\omega(x^{(k)})=0.$
Writing
$\omega(x^{(k)})=(1,\eta^{(k)\top})^{\top},\qquad\eta^{(k)}\in\mathbb{R}^{|\mathcal{F}|-1},$
the population vector $\eta^{(k)}$ solves
$H_{-}^{(k)}\eta^{(k)}=-h_{0}^{(k)}.$
When $\operatorname{rank}H^{(k)}=|\mathcal{F}|-1$, this system has the unique least-squares solution
$\eta^{(k)}=-\Big(H_{-}^{(k)\top}H_{-}^{(k)}\Big)^{-1}H_{-}^{(k)\top}h_{0}^{(k)}.$
We therefore define the sample estimator
$\hat{\eta}^{(k)}:=-\Big(\hat{H}_{-}^{(k)\top}\hat{H}_{-}^{(k)}\Big)^{-1}\hat{H}_{-}^{(k)\top}\hat{h}_{0}^{(k)},\qquad\hat{\omega}(x^{(k)}):=(1,\hat{\eta}^{(k)\top})^{\top}.$ (18)
Because the true normalized weight vector is strictly positive, positivity of $\hat{\omega}(x^{(k)})$ holds with probability approaching one.
Second-stage estimator: affine gate parameters.
For each non-baseline rule $f\neq f_{0}$, define
$\gamma_{f}:=(\alpha_{f},\beta_{f}^{\top})^{\top}\in\mathbb{R}^{d+1},\qquad z_{k}:=(1,x^{(k)\top})^{\top}.$
The affine gate implies
$\log\omega_{f}(x^{(k)})=z_{k}^{\top}\gamma_{f},\qquad k=1,\ldots,K.$
Hence, once the first-stage weights are estimated, the second stage is a linear regression of estimated log-weights on the identifying feature values:
$\hat{y}_{f}^{(k)}:=\log\hat{\omega}_{f}(x^{(k)}).$
Let
$X_{K}:=\begin{pmatrix}z_{1}^{\top}\\ \vdots\\ z_{K}^{\top}\end{pmatrix}$.
For any positive definite weight matrix $W_{f}\in\mathbb{R}^{K\times K}$, define the minimum-distance estimator
$\hat{\gamma}_{f}:=\operatorname*{arg\,min}_{\gamma\in\mathbb{R}^{d+1}}\big(\hat{y}_{f}-X_{K}\gamma\big)^{\top}W_{f}\big(\hat{y}_{f}-X_{K}\gamma\big)=(X_{K}^{\top}W_{f}X_{K})^{-1}X_{K}^{\top}W_{f}\hat{y}_{f}.$ (19)
When $K=d+1$, this collapses to exact inversion:
$\hat{\gamma}_{f}=X_{K}^{-1}\hat{y}_{f}.$
Population functionals.
The two-step estimator identifies the economically meaningful functionals of the gate. On the menu support $\mathcal{M}$, examples include:
$\bar{q}_{f}(\theta):=\frac{1}{M}\sum_{m=1}^{M}q_{f}(A_{m};\theta),$
$\bar{a}_{f}(\theta):=\frac{1}{M}\sum_{m=1}^{M}q_{f}(A_{m};\theta)\,A_{f}(A_{m}),$
$w_{f}(\theta):=\frac{1}{M}\sum_{m=1}^{M}\frac{q_{f}(A_{m};\theta)\,A_{f}(A_{m})}{\sum_{g\in\mathcal{F}}q_{g}(A_{m};\theta)\,A_{g}(A_{m})}.$
These correspond, respectively, to average latent gate mass, average effective active mass, and average conditional-on-activity responsibility.
Assumption 1.
The following hold:
  1. (E1)
    Interior probabilities. There exists $\underline{p}\in(0,1/2)$ such that $p(A_{m})\in[\underline{p},1-\underline{p}]$ for all $m$.
  2. (E2)
    Cellwise identification. For each $k=1,\ldots,K$, the matrix $H^{(k)}$ has rank $|\mathcal{F}|-1$.
  3. (E3)
    Affine richness. The matrix $X_{K}$ has full column rank $d+1$.
  4. (E4)
    Interior weights. For every $k$ and every $f\neq f_{0}$, $\omega_{f}(x^{(k)})\geq\underline{\omega}$ for some $\underline{\omega}>0$.
  5. (E5)
    Stable sampling shares. $n_{m}/N\to\pi_{m}\in(0,\infty)$ for each menu $m$.
Theorem 3 (Consistency and asymptotic normality of the two-step estimator).
Under Assumption 1:
  1. (i)
    For each $k=1,\ldots,K$,
    $\hat{\omega}(x^{(k)})\xrightarrow{p}\omega(x^{(k)}),\qquad\sqrt{N}\Big(\hat{\omega}(x^{(k)})-\omega(x^{(k)})\Big)\ \xrightarrow{d}\ \mathcal{N}\big(0,V_{\omega}^{(k)}\big).$
    In particular, applying the log transformation under (E4),
    $\sqrt{N}\Big(\hat{y}_{f}^{(k)}-\log\omega_{f}(x^{(k)})\Big)\ \xrightarrow{d}\ \mathcal{N}\big(0,V_{y,f}^{(k)}\big)\quad\text{for each }f\neq f_{0}.$
    Stacking across cells, $\sqrt{N}(\hat{y}_{f}-y_{f})\xrightarrow{d}\mathcal{N}(0,V_{y,f})$ where $\hat{y}_{f}:=(\hat{y}_{f}^{(1)},\ldots,\hat{y}_{f}^{(K)})^{\top}$ and $V_{y,f}=\operatorname{diag}(V_{y,f}^{(1)},\ldots,V_{y,f}^{(K)})$.
  2. (ii)
    For each non-baseline rule $f\neq f_{0}$,
    $\hat{\gamma}_{f}\xrightarrow{p}\gamma_{f},\qquad\sqrt{N}\big(\hat{\gamma}_{f}-\gamma_{f}\big)\ \xrightarrow{d}\ \mathcal{N}\big(0,V_{\gamma,f}\big),$
    where
    $V_{\gamma,f}=(X_{K}^{\top}W_{f}X_{K})^{-1}X_{K}^{\top}W_{f}\,V_{y,f}\,W_{f}X_{K}(X_{K}^{\top}W_{f}X_{K})^{-1}.$
  3. (iii)
    For any functional $\tau(\theta)$ that is continuously differentiable at the true parameter vector,
[…]
    is consistent and satisfies
    $\sqrt{N}\big(\hat{\tau}-\tau(\theta_{0})\big)\ \xrightarrow{d}\ \mathcal{N}\big(0,\nabla\tau(\theta_{0})^{\top}V_{\theta}\nabla\tau(\theta_{0})\big),$
    where $V_{\theta}$ is the asymptotic covariance matrix of the stacked estimator $\hat{\theta}=\{\hat{\gamma}_{f}\}_{f\neq f_{0}}$.
Proof sketch.
For each menu $m$, the sample share $\hat{p}_{m}$ is a sample mean of i.i.d. Bernoulli variables, so
$\sqrt{n_{m}}\big(\hat{p}_{m}-p(A_{m})\big)\xrightarrow{d}\mathcal{N}\big(0,p(A_{m})(1-p(A_{m}))\big).$
Because r(p)=p/(1p)r(p)=p/(1-p) is continuously differentiable on $\underline{p},1-\underline{p}]$, the delta method gives
$\sqrt{n_{m}}\big(\hat{r}_{m}-r(A_{m})\big)\xrightarrow{d}\mathcal{N}(0,\sigma_{r,m}^{2}).$
Hence each $\hat{H}^{(k)}$ is a smooth perturbation of $H^{(k)}$. Under (E2), $H_{-}^{(k)\top}H_{-}^{(k)}$ is nonsingular, so the map
$H^{(k)}\mapsto\eta^{(k)}=-\big(H_{-}^{(k)\top}H_{-}^{(k)}\big)^{-1}H_{-}^{(k)\top}h_{0}^{(k)}$
is continuously differentiable in a neighborhood of the truth. A second delta-method step yields consistency and asymptotic normality of $\hat{\eta}^{(k)}$, hence of $\hat{\omega}(x^{(k)})$ and $\hat{y}_{f}^{(k)}$. The second-stage estimator (19) is a linear map of $\hat{y}_{f}$, so its asymptotic properties follow immediately. The final statement is a standard delta-method consequence. The full proof is given in Appendix D. ∎
Efficient weighting and feasible inference.
A natural choice is feasible GLS, using a consistent estimator $\hat{V}_{y,f}$ of the first-stage covariance matrix and setting $W_{f}=\hat{V}_{y,f}^{-1}$. In practice, because the covariance formulas inherit the first-stage delta-method Jacobian, a cluster bootstrap at the menu level is attractive and directly yields standard errors for both the affine coefficients and the decomposition functionals $\bar{q}_{f}$, $\bar{a}_{f}$, and $w_{f}$.
Overidentification test of the affine gate.
When $K>d+1$, the affine gate restriction is overidentified. For each rule $f\neq f_{0}$, define the second-stage residual vector
$\hat{u}_{f}:=\hat{y}_{f}-X_{K}\hat{\gamma}_{f}.$
A natural minimum-distance specification test is
$J_{f}:=N\,\hat{u}_{f}^{\top}\hat{V}_{y,f}^{-1}\hat{u}_{f},$
where $\hat{V}_{y,f}$ is a consistent estimator of $V_{y,f}$ (the asymptotic variance of $\sqrt{N}(\hat{y}_{f}-y_{f})$ from Theorem 3). Under efficient weighting ($W_{f}=V_{y,f}^{-1}$) and the null that $\log\omega_{f}(x)$ is affine in $x$, standard minimum-distance theory gives the asymptotic reference distribution $J_{f}\xrightarrow{d}\chi^{2}_{K-d-1}$. A joint test across rules can be formed analogously by stacking the residuals.
The two-step estimator targets identification-consistent parameter recovery and supports inference on economically meaningful functionals. For out-of-sample prediction and model comparison, we rely on the MSE-based predictive estimator described in Section 6.5.
6 Data and empirical design
This section introduces the data and evaluation protocol. We describe the dataset, define the prediction target and evaluation metric, explain the gate-feature specification on which the identification conditions of Section 3 are verified, and detail the cross-validation protocol used for predictive comparison. Benchmark models are described at the end of the section.
@@ 6.1 Dataset and target variable
We use the choices13k dataset introduced by Peterson et al. (2021), containing 13,006 binary risky-choice problems. Each problem corresponds to a menu $A=\{L^{1},L^{2}\}$ and is repeated across many observations, allowing the computation of a menu-level choice frequency. We work with the filtered dataset that (i) removes ambiguous problems (masked probabilities) and (ii) keeps only problems for which feedback was provided, yielding T=9,831T=9{,}831 menus. The feedback restriction follows the baseline protocol of Peterson et al. (2021) and has a practical justification: feedback problems have more repetitions per menu, producing more precise estimates of the choice frequency $\hat{p}(A)$ that serves as our prediction target. Because our model is static (it maps menu features to a predicted choice probability without using feedback information), the feedback condition affects only the precision of the dependent variable, not the model’s information set. [^7] Feedback can in principle induce learning effects. If such dynamics are present, our static model captures the average behavior under the feedback regime. For each menu $A$, the dataset reports the fraction of times the left option is chosen. We denote this empirical frequency by $\hat{p}(A)$ and treat it as the target to be predicted.
@@ 6.2 Evaluation metric
Because the number of trials $n(A_{t})$ behind each menu-level frequency $\hat{p}(A_{t})$ varies across menus, we also report a trial-weighted variant:
$\mathrm{MSE}_{w}(g)=\frac{\sum_{t=1}^{T}n(A_{t})\Big(g(A_{t})-\hat{p}(A_{t})\Big)^{2}}{\sum_{t=1}^{T}n(A_{t})}.$ (21)
@@ 6.4 Cross-validation protocol
6.5 Predictive implementation
For the purpose of out-of-sample prediction and model comparison, we estimate the gate parameters by direct MSE minimization. This predictive estimator differs from the two-step econometric procedure of Section 5: the latter targets the identification-driven parameter recovery and supports inference for rule-level functionals, while the former is designed purely for predictive accuracy and benchmark evaluation. The rule-gating model differs from the benchmark neural predictors in a key way: the rules are deterministic and parameter-free. Estimation concerns only the gate, i.e. the mapping from menu features $z(A)$ to rule propensities $q(A)\in\Delta^{F-1}$ (see Section 6.6 for the definition of $z(A)$).
@@ 6.6 Gate features
The gate’s role is to modulate rule responsibilities using observable features of the menu. A natural question is which features to use. We adopt as the baseline specification a $12$-dimensional vector of economically motivated summary statistics $z(A)\in\mathbb{R}^{12}$. Six are pairwise gaps between the two lotteries: the expected-value gap $\mathrm{EV}(L^{1})-\mathrm{EV}(L^{2})$, the best-case outcome gap $\max(L^{1})-\max(L^{2})$, the worst-case outcome gap $\min(L^{1})-\min(L^{2})$, the variance gap $\mathrm{Var}(L^{1})-\mathrm{Var}(L^{2})$, the modal-outcome gap $\mathrm{Mode}(L^{1})-\mathrm{Mode}(L^{2})$, and the skewness gap $\mathrm{Skew}(L^{1})-\mathrm{Skew}(L^{2})$. These gaps are designed to capture the contrasts that the rules in the library operate on: the min/max gaps relate to MMn/MMx, the mode gap relates to MAP, and the EV and variance gaps relate to risk-return tradeoffs. The remaining six are per-lottery or menu-level features: the expected values $\mathrm{EV}(L^{1})$ and $\mathrm{EV}(L^{2})$, the standard deviations $\mathrm{SD}(L^{1})$ and $\mathrm{SD}(L^{2})$, the largest absolute outcome across both lotteries $\max_{i\in\{1,2\}}\max|\mathrm{Supp}(L^{i})|$, and the support-size gap $|\mathrm{Supp}(L^{1})|-|\mathrm{Supp}(L^{2})|$ (itself a pairwise difference, but included in this group because it measures menu complexity rather than a preference-relevant contrast). These features allow the gate to condition rule propensities on the scale and complexity of the menu environment. By construction, $d_{z}=12$ is much smaller than the $4M=40$ dimensions of the raw encoding, but it captures the economically relevant variation: each feature has a direct verbal interpretation as a lottery comparison or a menu characteristic.
@@
6.7 Benchmark models
$g(A;\theta)=\frac{\exp(\eta V(L^{1}))}{\exp(\eta V(L^{1}))+\exp(\eta V(L^{2}))},$ (23)
where $\eta$ is a free “determinism” (inverse-noise) parameter. The exception is the fully context-dependent network, which directly outputs $g(A;\theta)$.
Neural EU.
Neural PT.
Neural CPT.
Value-Based.
$V(L)=f(x^{L},p^{L}),$
Context-Dependent.
$g(A;\theta)=f_{\mathrm{NN}}(x^{L^{1}},p^{L^{1}},x^{L^{2}},p^{L^{2}}),$
Mixture of Theories (MOT).
7 Application: Procedural Heterogeneity in Risky Choice